OpenStudy (anonymous):

The temperature of a chemical reaction ranges between 40°C and 180°C. The temperature is at its lowest point when t = 0, the temperature is 90 degrees and increasing. What is a sine function that would model this reaction?

OpenStudy (anonymous):

Answer Choices: f(x) = 70 sin 12x + 110 f(x) = 110 sin 12x + 70 f(x) = 70 sin pi/6x + 110 f(x) = 110 sin pi/6x + 70

OpenStudy (anonymous):

I'm pretty sure it's D, but if someone could tell me if that's correct or not I'd really appreciate it.

OpenStudy (anonymous):

@ikram002p @jacquyjyz Would y'all mind taking a quick look at this when you get a chance?

OpenStudy (anonymous):

@ranga Could you please look at this when you get a chance?

OpenStudy (anonymous):

@dan815 Could you please help me?